# suppSupplementary References

### v2

@@ Acknowledgements
[…]
which is equivalent to the inequality
$\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau~{}\cdot~{}\left(\int_{\overline{A}}|f(\tau)|^{p-1}\,\mathrm{d}\tau+\int_{B\cup\overline{B}}|f(\tau)|^{p-1}\,\mathrm{d}\tau\right)\leq\int_{A}|f(\tau)|^{p-1}\,\mathrm{d}\tau~{}\cdot~{}\left(\int_{\overline{A}}|f(\tau)|^{p}\,\mathrm{d}\tau+\int_{B\cup\overline{B}}|f(\tau)|^{p}\,\mathrm{d}\tau\right).$
By adding the term $\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau\cdot\int_{A}|f(\tau)|^{p-1}\,\mathrm{d}\tau$ to both sides, we end up with the inequality
$\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau\int_{0}^{1}|f(\tau)|^{p-1}\,\mathrm{d}\tau\leq\int_{A}|f(\tau)|^{p-1}\,\mathrm{d}\tau\int_{0}^{1}|f(\tau)|^{p}\,\mathrm{d}\tau,$
@@
$\frac{\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau}{\int_{0}^{1}|f(\tau)|^{p}\,\mathrm{d}\tau}\leq\frac{\int_{A}|f(\tau)|^{p-1}\,\mathrm{d}\tau}{\int_{0}^{1}|f(\tau)|^{p-1}\,\mathrm{d}\tau}$
@@
$\frac{\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)}{\operatorname{WD}_{p}(F,G)}\leq\frac{\operatorname{Shift}_{+}^{\operatorname{WD}_{p-1}}(F,G)}{\operatorname{WD}_{p-1}(F,G)}.$
@@
00 0 $\leq\frac{\operatorname{CD}(F,G)}{\operatorname{Disp}_{+}^{\operatorname{CD}}(F,G)+\operatorname{Disp}_{-}^{\operatorname{CD}}(F,G)}-\frac{\operatorname{AVM}(F,G)}{\operatorname{Disp}_{+}^{\operatorname{AVM}}(F,G)+\operatorname{Disp}_{-}^{\operatorname{AVM}}(F,G)}$
$=\frac{2\widetilde{\rho}\phi(\widetilde{\mu}/\widetilde{\rho})+\widetilde{\mu}(2\Phi(\widetilde{\mu}/\widetilde{\rho})-1)-\sqrt{2}\phi(0)(\sigma_{F}+\sigma_{G})}{2\widetilde{\rho}\phi(0)-\sqrt{2}\phi(0)(\sigma_{F}+\sigma_{G})}-\frac{(2\Phi(\widetilde{\mu}/\widetilde{\sigma})-1)\widetilde{\mu}+2\widetilde{\sigma}\phi(\widetilde{\mu}/\widetilde{\sigma})}{2\widetilde{\sigma}\phi(0)}=:f(\widetilde{\mu}).$
@@
$f^{\prime}(\widetilde{\mu})$ $=\frac{2\widetilde{\rho}\phi^{\prime}(\widetilde{\mu}/\widetilde{\rho})/\widetilde{\rho}+\widetilde{\mu}(2\phi(\widetilde{\mu}/\widetilde{\rho})/\widetilde{\rho})+(2\Phi(\widetilde{\mu}/\widetilde{\rho})-1)}{2\widetilde{\rho}\phi(0)-\sqrt{2}\phi(0)(\sigma_{F}+\sigma_{G})}-\frac{(2\Phi(\widetilde{\mu}/\widetilde{\sigma})-1)+(2\phi(\widetilde{\mu}/\widetilde{\sigma})/\widetilde{\sigma})\widetilde{\mu}+2\widetilde{\sigma}\phi^{\prime}(\widetilde{\mu}/\widetilde{\sigma})/\widetilde{\sigma}}{2\widetilde{\sigma}\phi(0)}$
@@
Note that $f^{\prime}(0)=0$ and hence it suffices to show that the second derivative $f^{\prime\prime}(\widetilde{\mu})$ is non-negative for $\widetilde{\mu}>0$. The second derivative is given by
$f^{\prime\prime}(\widetilde{\mu})$ $=\frac{2\phi(\widetilde{\mu}/\widetilde{\rho})/\widetilde{\rho}}{2\widetilde{\rho}\phi(0)-\sqrt{2}\phi(0)(\sigma_{F}+\sigma_{G})}-\frac{2\phi(\widetilde{\mu}/\widetilde{\sigma})/\widetilde{\sigma}}{2\widetilde{\sigma}\phi(0)}$
@@
W.l.o.g. let $\sigma_{F}<\sigma_{G}=\sigma_{F}+\widetilde{\sigma}$. Note that $\widetilde{\rho}-\frac{1}{\sqrt{2}}(\sigma_{F}+\sigma_{G})=\sqrt{\sigma_{F}^{2}+(\sigma_{F}+\widetilde{\sigma})^{2}}-\frac{1}{\sqrt{2}}(2\sigma_{F}+\widetilde{\sigma})$ is zero if $\widetilde{\sigma}=0$ and has positive derivative
$\frac{\partial}{\partial\widetilde{\sigma}}(\sqrt{\sigma_{F}^{2}+(\sigma_{F}+\widetilde{\sigma})^{2}}-\frac{1}{\sqrt{2}}(2\sigma_{F}+\widetilde{\sigma}))$ $=\frac{\sigma_{F}+\widetilde{\sigma}}{\sqrt{\sigma_{F}^{2}+(\sigma_{F}+\widetilde{\sigma})^{2}}}-\frac{1}{\sqrt{2}}=\frac{1}{\sqrt{\sigma_{F}^{2}/(\sigma_{F}+\widetilde{\sigma})^{2}+1}}-\frac{1}{\sqrt{2}}$
if $\widetilde{\sigma}>0$. Hence, the denominator in expression (S8) is positive. Therefore, it suffices to show that the numerator in expression (S8) is also non-negative. As $\widetilde{\rho}=\sqrt{\sigma_{F}^{2}+\sigma_{G}^{2}}=\sqrt{(\sigma_{F}-\sigma_{G})^{2}+2\sigma_{F}\sigma_{G}}>\sqrt{(\sigma_{F}-\sigma_{G})^{2}}=\widetilde{\sigma}$ and hence $\phi(\widetilde{\mu}/\widetilde{\rho})>\phi(\widetilde{\mu}/\widetilde{\sigma})$, it suffices to show that $\widetilde{\sigma}/\widetilde{\rho}\geq(\widetilde{\rho}-\frac{1}{\sqrt{2}}(\sigma_{F}+\sigma_{G}))/\widetilde{\sigma}$ or, equivalently,
11 11 $\geq\widetilde{\rho}(\widetilde{\rho}-\frac{1}{\sqrt{2}}(\sigma_{F}+\sigma_{G}))/\widetilde{\sigma}^{2}$
$=\sqrt{(\sigma_{F}/\widetilde{\sigma})^{2}+(\sigma_{F}/\widetilde{\sigma}+1)^{2}}\left(\sqrt{(\sigma_{F}/\widetilde{\sigma})^{2}+(\sigma_{F}/\widetilde{\sigma}+1)^{2}}-\left(\sigma_{F}/\widetilde{\sigma}+(\sigma_{F}/\widetilde{\sigma}+1)\right)/\sqrt{2}\right)=:g(\sigma_{F}/\widetilde{\sigma}).$
@@
for $y\geq 0$ (note that the numerator of the fraction under the root in the last term is clearly larger than the denominator), $g(x)=h(2x^{2}+2x)$ is also decreasing in $x$ and, hence, the inequality $1>g(0)>g(\sigma_{F}/\widetilde{\sigma})$ holds, which finishes the proof of the theorem. ∎
@@
    If $F>_{\text{D}}G$, then there exist $10<\xi<\tau<$ 1 such that the defining inequality
@@
      $-f(\xi)>\underset{\mathrlap{=0}}{\underbrace{g(\tau)-f(\tau)}}-g(\xi)\quad\Longrightarrow\quad f(1-\xi)>g(1-\xi).$
@@
[…]
@@
Clearly, the weak stochastic order is reflexive. To show the transitivity, consider three distributions $F\geq_{\text{wS}}G\geq_{\text{wS}}H$, and arbitrary $10.5<\tau,\xi<$ 1, which are fixed throughout the proof. The distribution $F$ is larger than $H$ in weak stochastic order if the following inequality is satisfied
00 0 $\leq\max\{F^{-1}(\tau)-H^{-1}(\xi),F^{-1}(1-\tau)-H^{-1}(1-\xi)\}$
@@
where $10.5<\gamma<$ 1. For any $\gamma$, at least one of the terms $C(\gamma)$ or $D(\gamma)$ is non-negative by definition of the weak stochastic ordering $G\geq_{\text{wS}}H$. Therefore, it suffices to show that there exists a $\gamma_{0}\in[0.5,1]$ such that both $A(\gamma_{0})$ and $B(\gamma_{0})$ are non-negative to prove the above inequality. To this end, let $\gamma_{0}=\sup\{\gamma\mid A(\gamma)\geq 0\}$. Note that $\gamma_{0}\geq 0.5$ as $F^{-1}(\tau)\geq G^{-1}(0.5)$ by weak stochastic ordering (which, by setting $\xi=0.5$ in its definition, implies that either $F^{-1}(\tau)\geq G^{-1}(0.5)$ or $F^{-1}(1-\tau)\geq G^{-1}(0.5)$ and $F^{-1}(\tau)\geq F^{-1}(1-\tau)$ holds as $\tau>0.5$.)
@@
HadGEM2-CC Met Office Hadley Centre, UK

### v3

@@ Acknowledgements
$\leq\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau$ $\cdot$~ $\bigg{(}\int_{A}|f(1-\tau)|^{q}\,\mathrm{d}\tau+\int_{B}|f(\tau)|^{q}+|f(1-\tau)|^{q}\,\mathrm{d}\tau\bigg{)},$
which is equivalent (as $\tau\in A\Longleftrightarrow(1-\tau)\in\overline{A}$, and analogously for $B$ and $\overline{B}$) to the inequality
$\int_{A}|f(\tau)|^{q}\,\mathrm{d}\tau~\cdot~\left(\int_{\overline{A}}|f(\tau)|^{p}\,\mathrm{d}\tau+\int_{B\cup\overline{B}}|f(\tau)|^{p}\,\mathrm{d}\tau\right)\leq\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau~\cdot~\left(\int_{\overline{A}}|f(\tau)|^{q}\,\mathrm{d}\tau+\int_{B\cup\overline{B}}|f(\tau)|^{q}\,\mathrm{d}\tau\right).$
By adding the term $\int_{A}|f(\tau)|^{q}\,\mathrm{d}\tau\cdot\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau$ to both sides, we end up with the inequality
$\int_{A}|f(\tau)|^{q}\,\mathrm{d}\tau\int_{0}^{1}|f(\tau)|^{p}\,\mathrm{d}\tau\leq\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau\int_{0}^{1}|f(\tau)|^{q}\,\mathrm{d}\tau,$
@@
$\frac{\int_{A}|f(\tau)|^{q}\,\mathrm{d}\tau}{\int_{0}^{1}|f(\tau)|^{q}\,\mathrm{d}\tau}\leq\frac{\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau}{\int_{0}^{1}|f(\tau)|^{p}\,\mathrm{d}\tau}$
@@
$\frac{\operatorname{Shift}_{+}^{\operatorname{WD}_{q}}(F,G)}{\operatorname{WD}_{q}(F,G)}\leq\frac{\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)}{\operatorname{WD}_{p}(F,G)}.$
@@
$0$ $\leq\frac{\operatorname{CD}(F,G)}{\operatorname{Disp}_{+}^{\operatorname{CD}}(F,G)+\operatorname{Disp}_{-}^{\operatorname{CD}}(F,G)}-\frac{\operatorname{AVM}(F,G)}{\operatorname{Disp}_{+}^{\operatorname{AVM}}(F,G)+\operatorname{Disp}_{-}^{\operatorname{AVM}}(F,G)}$
[…]
@@
[…]
@@
[…]
@@
[…]
$)\displaystyle\frac{\partial}{\partial\widetilde{\sigma}}(\sqrt{\sigma_{F}^{2}+(\sigma_{F}+\widetilde{\sigma})^{2}}-\frac{1}{\sqrt{2}}(2\sigma_{F}+\widetilde{\sigma}$)) $=\frac{\sigma_{F}+\widetilde{\sigma}}{\sqrt{\sigma_{F}^{2}+(\sigma_{F}+\widetilde{\sigma})^{2}}}-\frac{1}{\sqrt{2}}=\frac{1}{\sqrt{\sigma_{F}^{2}/(\sigma_{F}+\widetilde{\sigma})^{2}+1}}-\frac{1}{\sqrt{2}}$
[…]
$1$ $\geq\widetilde{\rho}(\widetilde{\rho}-\frac{1}{\sqrt{2}}(\sigma_{F}+\sigma_{G}))/\widetilde{\sigma}^{2}$
[…]
@@
[…]
@@
    If $F>_{\text{D}}G$, then there exist $0<\xi<\tau<1$ such that the defining inequality
@@
      $f(\xi)>\underset{\mathrlap{=0}}{\underbrace{g(\tau)-f(\tau)}}-g(\xi)\quad\Longrightarrow\quad f(1-\xi)>g(1-\xi).$
@@
    We prove the reverse implication by contraposition. If $F\not\geq_{\text{wS}}G$, then there exists a pair $(\tau,\xi)\in(\frac{1}{2},1)^{2}$ such that $\max\big{\{}F^{-1}(\tau)-G^{-1}(\xi),\,F^{-1}(1-\tau)-G^{-1}(1-\xi)\big{\}}\;<\;0$. By continuity of the quantile functions there is a neighborhood $B\subset(\frac{1}{2},1)^{2}$ of $(\tau,\xi)$ such that the above inequality holds for all pairs $(\tau,\xi)\in B$. Substitution with $\tau=\frac{1+\beta}{2}$ and $\xi=\frac{1+\alpha}{2}$ in the integral for the shift component of the $\operatorname{CD}$ (as in (16) with $F$ and $G$ switched) yields
@@
Clearly, the weak stochastic order is reflexive. To show the transitivity, consider three distributions $F\geq_{\text{wS}}G\geq_{\text{wS}}H$, and arbitrary $0.5<\tau,\xi<1$, which are fixed throughout the proof. The distribution $F$ is larger than $H$ in weak stochastic order if the following inequality is satisfied
$0$ $\leq\max\{F^{-1}(\tau)-H^{-1}(\xi),F^{-1}(1-\tau)-H^{-1}(1-\xi)\}$
@@
where $0.5<\gamma<1$. For any $\gamma$, at least one of the terms $C(\gamma)$ or $D(\gamma)$ is non-negative by definition of the weak stochastic ordering $G\geq_{\text{wS}}H$. Therefore, it suffices to show that there exists a $\gamma_{0}\in[0.5,1]$ such that both $A(\gamma_{0})$ and $B(\gamma_{0})$ are non-negative to prove the above inequality. To this end, let $\gamma_{0}=\sup\{\gamma\mid A(\gamma)\geq 0\}$. Note that $\gamma_{0}\geq 0.5$ as $F^{-1}(\tau)\geq G^{-1}(0.5)$ by weak stochastic ordering (which, by setting $\xi=0.5$ in its definition, implies that either $F^{-1}(\tau)\geq G^{-1}(0.5)$ or $F^{-1}(1-\tau)\geq G^{-1}(0.5)$ and $F^{-1}(\tau)\geq F^{-1}(1-\tau)$ holds as $\tau>0.5$.)
@@
HadGEM2-$C$ Met Office Hadley Centre, UK